OpenStudy (anonymous):

If xy + 8e^y = 8e, find the value of y'' at x =0. Attempt: I know what the second, and first derivative will look like. I just don't know how to go about with plugging in the x=0. If xy + 8e^y = 8e, find the value of y'' at x =0. Attempt: I know what the second, and first derivative will look like. I just don't know how to go about with plugging in the x=0. @Mathematics

OpenStudy (anonymous):

@amytincan what r u getting the value of y''? After knowing the value of y'' , it will be possible to plug the value of x

OpenStudy (anonymous):

I'm not doing so well on this I'm at y' \[y'=-y/(x+8e^yln8e)\]

OpenStudy (anonymous):

i m getting y' = -y/x+8e^y

OpenStudy (anonymous):

Niksva, I am stuck on the denominator, apparently I do not know how to get the derivative of 8e^y. Is the formula for this portion supposed to be d/dx[a^g(x)]=a^g(x)lna(g'x)?

OpenStudy (anonymous):

no for the exponential function we follow the formula given below d/dp (e^q) = e^q dq/dp we write the exponential function as it is

OpenStudy (anonymous):

But that's a particular case of @amytincan 's Formula... when a=e @niksva ... Because lne =1 :-)

OpenStudy (anonymous):

omg my fault sorry @amytincan and thanx @salonigupta95 but i have done the first derivative right

OpenStudy (anonymous):

Yeah... I agree in that ...

OpenStudy (anonymous):

it seems second derivative is going to get more complex

OpenStudy (anonymous):

Is it fine???  Verify once...